How many grams of silver bromide, AgBr, are produced if 15.0 mL of 0.350 M CaBr2(aq) completely react given the reaction:

2 AgNO3(aq) + CaBr2(aq) → Ca(NO3)2(aq) + 2 AgBr(s)
1
Expert's answer
2019-04-29T03:15:26-0400

n(CaBr2)= V(CaBr2)×c(CaBr2)=0.015×0.35=0.00525 mol

n(AgBr)=2×n(CaBr2)=2×0.00525=0.0105 mol

m(AgBr)=n(AgBr)×M(AgBr)=0.0105×188=1.974 g.